# Efficient Multimodal Streaming Recommendation via Expandable Side Mixture-of-Experts

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 4.3. Expert Pruning
where $\|\cdot\|_{2}$ is the Euclidean norm and NN is the number of experts at that layer. Note that the subscript indexing the layers are omitted for simplicity. The numerator measures the magnitude of expert $E_{j}$’s contribution to the final output, while the denominator aggregates the total contribution from all expert sources, including the frozen backbone and other experts. As a result, $r_{j}\in[0,1]$ quantifies the relative influence of expert $E_{j}$ in producing the layer’s output.
@@
$4.$ Sequence Encoding and Loss Function
@@ 4.5. Complexity Analysis
Hereby we analyze the computational complexity of XSMoE w.r.t. parameter efficiency, training time complexity, and GPU memory efficiency per modality, following the protocols in (Fu et al., 2024a). We denote the input and output size of each expert as $d$, the number of layers in each side-tuning network as $M$, the down-projection size as $d^{\prime}$, and the number of experts at each layer as NN. Typically, we have $N<d^{\prime}<d$ due to the expert pruning mechanism. For simplicity, we assume each layer contains the same number of experts NN.
@@ 4.5.1. Parameter Efficiency
Each layer of XSMoE contains two components: NN FFN experts and 1 router. Each expert contains one up-projection layer and one down-projection layer, both of which has $dd^{\prime}$ parameters. Therefore, the parameter size of an expert is $2dd^{\prime}$. The input size of the router is $d$ and the output size is $N+1$. Hence, each router has $dN+d$ parameters. Therefore, given $M$ layers, each with NN experts and one router, the number of parameters of XSMoE is $M(2Ndd^{\prime}+Nd+d)\approx\mathcal{O}(MNdd^{\prime})$.
@@ 4.5.3. GPU Memory Complexity
The GPU memory usage is dominated by (1) model weights, (2) gradients, (3) optimizer states, and (4) activations (Fu et al., 2024a). At each layer, (1) is $\mathcal{O}(Ndd^{\prime})$. (2) only involves trainable parameters and is thus $\mathcal{O}(dd^{\prime})$. The optimizer states of Adam double the trainable parameter count are thus $\mathcal{O}(2dd^{\prime})$. The activations are stored after forward passes and are used for backpropagation. In XSMoE, activations are stored for the router and the experts. At each layer, the router that computes Softmax scores NN experts and its corresponding backbone encoder. Hence, the activations for the routers are $\mathcal{O}(N+1)$. For NN experts, the activations are $\mathcal{O}(Nd)$. In summary, the GPU memory complexity of XSMoE is $\mathcal{O}(Ndd^{\prime})+\mathcal{O}(dd^{\prime})+\mathcal{O}(2dd^{\prime})+\mathcal{O}(N+1)+\mathcal{O}(Nd)\approx\mathcal{O}(Ndd^{\prime})$ for each layer and $\mathcal{O}(MNdd^{\prime})$ for $M$ layers.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; November 10–14, 2025; Seoul, Koreaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 4.3. Expert Pruning
where $\|\cdot\|_{2}$ is the Euclidean norm and $N$ is the number of experts at that layer. Note that the subscript indexing the layers are omitted for simplicity. The numerator measures the magnitude of expert $E_{j}$’s contribution to the final output, while the denominator aggregates the total contribution from all expert sources, including the frozen backbone and other experts. As a result, $r_{j}\in[0,1]$ quantifies the relative influence of expert $E_{j}$ in producing the layer’s output.
@@
4.4. Sequence Encoding and Loss Function
@@ 4.5. Complexity Analysis
Hereby we analyze the computational complexity of XSMoE w.r.t. parameter efficiency, training time complexity, and GPU memory efficiency per modality, following the protocols in (Fu et al., 2024a). We denote the input and output size of each expert as $d$, the number of layers in each side-tuning network as $M$, the down-projection size as $d^{\prime}$, and the number of experts at each layer as $N$. Typically, we have $N<d^{\prime}<d$ due to the expert pruning mechanism. For simplicity, we assume each layer contains the same number of experts $N$.
@@ 4.5.1. Parameter Efficiency
Each layer of XSMoE contains two components: $N$ FFN experts and 1 router. Each expert contains one up-projection layer and one down-projection layer, both of which has $dd^{\prime}$ parameters. Therefore, the parameter size of an expert is $2dd^{\prime}$. The input size of the router is $d$ and the output size is $N+1$. Hence, each router has $dN+d$ parameters. Therefore, given $M$ layers, each with $N$ experts and one router, the number of parameters of XSMoE is $M(2Ndd^{\prime}+Nd+d)\approx\mathcal{O}(MNdd^{\prime})$.
@@ 4.5.3. GPU Memory Complexity
The GPU memory usage is dominated by (1) model weights, (2) gradients, (3) optimizer states, and (4) activations (Fu et al., 2024a). At each layer, (1) is $\mathcal{O}(Ndd^{\prime})$. (2) only involves trainable parameters and is thus $\mathcal{O}(dd^{\prime})$. The optimizer states of Adam double the trainable parameter count are thus $\mathcal{O}(2dd^{\prime})$. The activations are stored after forward passes and are used for backpropagation. In XSMoE, activations are stored for the router and the experts. At each layer, the router that computes Softmax scores $N$ experts and its corresponding backbone encoder. Hence, the activations for the routers are $\mathcal{O}(N+1)$. For $N$ experts, the activations are $\mathcal{O}(Nd)$. In summary, the GPU memory complexity of XSMoE is $\mathcal{O}(Ndd^{\prime})+\mathcal{O}(dd^{\prime})+\mathcal{O}(2dd^{\prime})+\mathcal{O}(N+1)+\mathcal{O}(Nd)\approx\mathcal{O}(Ndd^{\prime})$ for each layer and $\mathcal{O}(MNdd^{\prime})$ for $M$ layers.